OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -6, an = 4 • an-1 PLEASE HELP

OpenStudy (anonymous):

so for a2 you plug 2 instead of n in the expression for an : a2 = 4* (a(2-1)) = 4*a1 = a1 is given = 4*-6 = -24

OpenStudy (anonymous):

show me how you find a3 now

OpenStudy (anonymous):

a3=4*(a(3-1))=4*a2=4*-24=-96?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

as you can see the next term will be 4*a3 etc

OpenStudy (anonymous):

THANK YOU!!!:)

OpenStudy (anonymous):

yw